# TSCMamba: Mamba Meets Multi-View Learning for Time Series Classification

### v1

@@ Abstract
Time series classification (TSC) on multivariate time series is a critical problem. We propose a novel multi-view approach integrating frequency-domain and time-domain features to provide complementary contexts for TSC. Our method fuses continuous wavelet transform spectral features with temporal convolutional or multilayer perceptron features. We leverage the Mamba state space model for efficient and scalable sequence modeling. We also introduce a novel tango scanning scheme to better model sequence relationships. Experiments on 10 standard benchmark datasets demonstrate our approach achieves an average 6.45% accuracy improvement over state-of-the-art TSC models.
@@ 1 Introduction
Time series classification (TSC) is a fundamental task in diverse fields abundant with time series data. With the advancing of sensing technologies, multivariate time series (MTS) data have been ubiquitous, and thus TSC over MTS has attracted ever-increasing research attention.
MTS data is usually highly redundant, with significant inter-observation correlations in temporal neighborhoods. This redundancy is similar to that found in images (Bengio et al., 2013), where convolutional kernels are effective in extracting features. Convolutional kernels can be used in the time-domain to extract features from MTS data, but they are limited by their small receptive field, thus typically extracting temporally localized contents. They have limited ability to capture spectrally localized contents or global interactions in the MTS. To remedy this limitation, frequency-domain features can be extracted using transforms such as the Fourier transform, digital wavelet transforms, and the shapelet transform. These transforms can help capture global interactions and spectral patterns in the data and separate noise from signals or useful contents.
MTS data is often noisy (Kang et al., 2014), which may include Gaussian noise, speckle noise, seasonal noise, trend noise, and outliers. Noise can mask the underlying signals, make classification more difficult, and affect the robustness of the algorithms. Thus, it is important to obtain effective representations that are robust to noise. While Transformers can be effective for capturing global interactions (Vaswani et al., 2017), they have been found to be sensitive to noise in the data (Mahmood et al., 2021).
Existing methods for TSC typically focus on a single type of features, such as temporal-domain features or patterns, or frequency-domain patterns. However, these methods do not fully exploit the complementary characteristics that can be obtained from different domains. For example, time-domain features may readily capture temporally local patterns, while frequency-domain features may help capture spectrally local and sensible patterns. Multi-view learning has been successful in capturing complementary characteristics from complex data, robust to noise or disturbances, and has been widely used in image classification and clustering (Peng et al., 2024). However, effective multi-view strategies have not been well investigated for enhancing deep learning-based TSC in the literature.
To address this gap, we propose a novel approach for TSC that effectively integrates features and patterns from both time- and frequency domains. For time-domain features, we leverage an efficient convolutional kernel-based feature representation or project patterns using a fully connected neural network. For frequency-domain patterns, we adopt features obtained from frequency-domain transformations, such as continuous wavelet transform (CWT). By integrating these features and patterns, we provide comprehensive multi-view contexts for TSC.
Convolutional kernels used in convolutional neural networks (CNNs) excel at capturing temporal dependencies between input features or inter-dependencies between channels. However, these kernels typically have limited lengths, resulting in a limited receptive field that captures local patterns. In contrast, multi-layer perceptrons (MLPs) are fully connected neural networks with a global receptive field, allowing them to learn global patterns and dependencies among input features. While MLPs can capture global patterns, they may not be well-suited for capturing temporally local patterns or temporal/spatial dependencies, as they treat each input feature independently and do not explicitly consider spatial relationships among neighboring features.
To capture spectral-domain patterns, we leverage CWT to decompose and represent a time series. CWT is an invertible wavelet transformation that can perfectly reconstruct the original data, preserving all information. This representation will provide a rich spectral view to complement the temporal views. While CWT has been used to extract features for fault diagnosis in the literature (Wang et al., 2020), it has not been used to develop a general model for TSC. To reduce the computational complexity of CWT, we will compute the CWT coefficients for each time series and use them as input for the spectral view.
The MTS data often has sensible global patterns, such as trend, seasonality, periodicity, cycles, and long-term dependencies. MTS may also comprise informational local patterns, such as peaks, troughs, jumps, local cycles, and plateaus with local trends. These global or local patterns may constitute key information to distinguish the classes of MTS. As different MTS datasets may have different global or local contents, it is critical to leverage sensible MTS patterns for TSC. Thanks to its multi-scale decomposition, the CWT representation for the spectral view contains global and local clues at various levels of the MTS. However, temporal views obtained from the CNN-based feature extraction and MLP-based feature maps may not provide global patterns or local cues as proper contexts for inference of class labels.
To effectively capture discriminative patterns in the temporal domain, we propose a switch mechanism that selects between CNN-based local features and MLP-based global patterns. We utilize the kernel-based feature transformation ROCKET (Dempster et al., 2020) to extract locally representative features while using a fully connected MLP to capture global patterns. The switch mechanism determines which temporal view (local or global) is more informative and integrates it with the CWT-domain features.
Consequently, we have diverse views, representing global temporal features, localized temporal features, and multi-scale spectral features. We will use them to provide discriminative and complementary contextual clues to an effective inference engine to capture long-term dependencies between these features. After obtaining features with these different views for time series data, we fuse them to provide rich contexts for subsequent sequence modeling. As discussed above, MLP is good at capturing global, but not local interactions. CNN can extract localized features but its receptive field is typically limited. We will use a switch gate to determine whether to fuse transform domain features with local features or global patterns in the temporal domain, which will be determined by tuning the model to select between these two approaches.
Having obtained salient contextual cues, we will use a state-of-the-art (SOTA) state-space model (SSM) driven inference engine, Mamba (Gu & Dao, 2023), to decide on the class labels. Mamba is an architecture for sequence modeling, which is a special case of SSM and similar to recursive neural networks (RNNs). The SSM is a classical concept in the fields of control and signal processing (Gu et al., 2021), which uses state variables to represent the system’s internal condition and describes how the state variables change over time based on the inputs and the current state. By introducing a new selective updating mechanism of the hidden states, called selective state spaces, it can selectively update a subset of state dimensions based on the input at each time step. As a result, Mamba can efficiently capture long-range dependencies. Mamba-based models have demonstrated competitive performance on various tasks, such as language modeling (Gu & Dao, 2023; Dao & Gu, 2024), time series forecasting (Ahamed & Cheng, 2024b), DNA sequence modeling (Gu & Dao, 2023), tabular data learning (Ahamed & Cheng, 2024a), and audio generation (Shams et al., 2024).
Unlike popular Transformers, which are sequence modeling architectures with quadratic time complexity in sequence length, Mamba achieves linear time complexity. This makes it more suitable for processing long sequences and scaling to larger datasets. Therefore, by using Mamba, our TSC model is efficient in training and inference, with reduced computational costs and memory requirements compared to existing SOTA models. Moreover, we introduce a novel sequence scanning scheme for the Mamba block, which uses essentially the same memory footprint but demonstrates higher accuracy than the vanilla Mamba scanning. Through extensive experiments, we will demonstrate that our multi-view Mamba-based approach outperforms or matches the performance of existing SOTA models that typically have more computationally expensive models, with a superior average accuracy over a range of 10 standard benchmarking datasets.
In summary, the contribution of this paper includes but is not limited to the following:
@@
  • We propose a novel multi-view approach for TSC, which seamlessly integrates frequency-domain and time-domain features to provide complementary and discriminative contexts for classification. In particular, we propose to fuse features from both domains with a gating scheme. Our approach can effectively leverage local and global patterns that characterize the MTS classes, enhancing the discriminative power.
  • We leverage a concurrent SSM technique, Mamba, for sequence modeling to capture long-term dependencies within the MTS with linear efficiency and scalability. Moreover, we propose an innovative Mamba-based scanning scheme, called tango scanning, to scan the integrated features to identify salient contents within complex contexts. This scanning is demonstrated to be more effective in modeling the relationships in the sequence than that of vanilla Mamba block for TSC.
  • We first use the CWT for multi-scale representation of MTS in the general task of TSC. This joint temporal and spectral transformation allows for efficient frequency-domain feature extraction and subsequent fusion of time-domain features with frequency-domain features.
@@
  • Our extensive experiments validate the proposed approach, which demonstrates superior performance to various existing SOTA models over 10 standard benchmarking datasets. On average, an improvement of 6.45% in accuracy is obtained.
With these contributions, we expect our novel approach to help advance real-world TSC applications in diverse fields of research and everyday life. In the following sections, we first briefly review related works, then presents our approach in detail, and finally demonstrate extensive experimental results and ablation studies to conclude the paper.
@@ 2 Related Works
Traditional TSC methods include techniques like Dynamic Time Warping (DTW) (Berndt & Clifford, 1994), which measures the similarity between time series by aligning them in a non-linear way of dynamic programming. Tree-based methods like XGBoost (Chen & Guestrin, 2016) have also been applied to the TSC task.
@@
In recent years, deep learning approaches have become increasingly popular for TSC. Various MLP-based methods have been proposed, including DLinear by Zeng et al. (2023) and LightTS by Zhang et al. (2022). DLinear constructs a simple model based on MLP, while LightTS uses light sampling-oriented MLP. These models are generally efficient in computations. Convolutional neural networks (CNNs) have been adapted for TSC, such as ROCKET (Dempster et al., 2020) which uses random convolutional kernels for fast and accurate classification. CNN has been also used by Franceschi et al. (2019) for learning representations of multivariate time series in an unsupervised way, which is then further leveraged for TSC. Besides CNNs, recursive neural networks (RNNs) such as long-short-term memory (LSTM) (Hochreiter & Schmidhuber, 1997) and a variant, a gated recursive unit (GRU), has also been adopted for TSC. Moreover, CNN and RNN have been combined to handle TSC effectively (Lai et al., 2018).
@@
Transformers by Vaswani et al. (2017), originally used for natural language processing, have been adapted for time series modeling. Reformer by Kitaev et al. (2020) introduces efficiency improvements to handle longer sequences. Numerous Transformer variants have been proposed to better model the unique characteristics of time series, such as handling non-stationarity with on-stationary Transformers by Liu et al. (2022), combining exponential smoothing with ETSformer in Woo et al. (2022), and using decomposition and auto-correlation with Autoformer in Wu et al. (2021). Other variants include Pyraformer (Liu et al., 2021), which uses pyramidal attention to reduce complexity, Flowformer (Wu et al., 2022b), which linearizes Transformers using conservation flows, and Informer (Zhou et al., 2022), which focuses on efficient long sequence forecasting by exploiting frequency-enhanced decomposition. Notably, TimesNet (Wu et al., 2022a) models the temporal 2D variations in time series data using a hierarchical structure of temporal blocks. By combining 2D convolutions, multi-head self-attention, and a novel positional encoding scheme, it can capture both local patterns and long-range dependencies in time series and obtain state-of-the-art performance. Despite the impressive performance of Transformer-based models, Zeng et al. (2023) have shown that MLP-based models can be more effective in many scenarios.
Recently, a method called LSSL by Gu et al. (2022) has been proposed for TSC with a structured SSM called S4. It employs a special parameterization called the Diagonal Plus Low-Rank form to represent the state transition matrix, enabling efficient computation over long sequences.

### v2

@@ Abstract
Multivariate time series classification (TSC) is critical for various applications in fields such as healthcare and finance. While various approaches for TSC have been explored, important properties of time series, such as shift equivariance and inversion invariance, are largely underexplored by existing works. To fill this gap, we propose a novel multi-view approach to capture patterns with properties like shift equivariance. Our method integrates diverse features, including spectral, temporal, local, and global features, to obtain rich, complementary contexts for TSC. We use continuous wavelet transform to capture time-frequency features that remain consistent even when the input is shifted in time. These features are fused with temporal convolutional or multilayer perceptron features to provide complex local and global contextual information. We utilize the Mamba state space model for efficient and scalable sequence modeling and capturing long-range dependencies in time series. Moreover, we introduce a new scanning scheme for Mamba, called tango scanning, to effectively model sequence relationships and leverage inversion invariance, thereby enhancing our model’s generalization and robustness. Experiments on two sets of benchmark datasets (10+20 datasets) demonstrate our approach’s effectiveness, achieving average accuracy improvements of 4.01-6.45% and 7.93% respectively, over leading TSC models such as TimesNet and TSLANet.
keywords:
Time Series Classification, Deep Learning , state-space-machine , multi-view learning
journal: Information Fusion
[…]
[cs]organization=Department of Computer Science, University of Kentucky, city=Lexington, state=KY, country=USA
[…]
[ibi]organization=Institute for Biomedical Informatics, University of Kentucky, city=Lexington, state=KY, country=USA
@@ 1 Introduction
Time series classification (TSC) is a fundamental task in diverse fields abundant with time series data such as healthcare, weather forecasting, and finance. With the advancement of sensing technologies, multivariate time series (MTS) data have been ubiquitous, and thus TSC over MTS has attracted increasing research attention. Various approaches for TSC have been explored, including statistical, signal processing, and machine or deep learning approaches in both time- and frequency domains. Despite intensive studies and rapid progress, important properties of many time series, such as shift equivariance and inversion invariance, remain underexplored in TSC. Moreover, existing methods face significant challenges in handling long-range dependencies, effectively utilizing both time and frequency representations, and maintaining computational efficiency. Many deep learning models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), primarily focus on time-domain features, often overlooking frequency-based representations that could provide additional discriminatory power. Transformer-based approaches, while powerful, suffer from quadratic complexity in sequence length, making them less efficient for long time-series data. Meanwhile, state-space models (SSMs) like Mamba offer promising alternatives but have not been widely explored in TSC. These challenges, including the need for better handling of time series properties and computational efficiency, motivate our work in developing a unified, efficient, and multi-view TSC approach leveraging both spectral and temporal insights.
Equivariance refers to a property where transformations applied to the input lead to predictable and corresponding transformations in the output. In the case of shift equivariance for time series data, this means that when we apply a shift transformation to the input time series, the output of our model (such as detected features or patterns) undergoes a corresponding shift. Shift equivariance property is crucial for TSC as it allows pattern recognition regardless of exact temporal positions, offering benefits such as:
  • 1.
    Resilience to temporal misalignments between different instances of the same class, which is common in real-world data.
  • 2.
    Creation of more consistent features across samples when dealing with MTS of varying lengths, such as when training and test examples have different durations.
  • 3.
    Improved generalization while retaining information about temporal relationships, as the classifier can learn patterns that are consistent in their relative positions across different time shifts.
These benefits collectively enhance the robustness and flexibility of TSC models, making them more adaptable to the diverse and often noisy nature of real-world time series data. However, existing approaches often struggle to fully exploit these advantages. While traditional statistical methods like dynamic time warping (DTW) excel at alignment, they struggle with complex temporal structures. Deep learning approaches each have their limitations: CNNs, though inherently shift-equivariant, struggle with long-range dependencies; Transformers offer flexible attention but at high computational cost; and SSMs like Mamba, while promising efficient sequence modeling, require specialized adaptations for TSC tasks. These limitations highlight the need for a TSC model that effectively balances efficiency, feature diversity, and sequence modeling capabilities.
Existing machine learning approaches for TSC often rely on convolution-based methods or neural networks to endow shift equivariance for the extracted features [1]. Many models have been designed to exploit this property for visual learning tasks with 2D or 3D data [2, 3]. For TSC, several convolution-based methods [4, 5, 6] have implicitly utilized this property. However, they only employ temporal features by applying convolutions in the time domain.
In addition to time-based features, frequency-based features can help identify important patterns in the data’s spectral composition. Disentangling time and frequency components has been shown to be critical for time series learning [7, 8]. In particular, spectral features have been extracted using discrete Fourier transform (DFT) [8] and discrete wavelet transform (DWT) [9, 10, 11]. However, neither DFT nor DWT is shift equivariant - DFT lacks this property [12], while DWT’s discrete nature and downsampling (decimation) step in its computation also prevent shift equivariance [13, 14]. Consequently, a small shift in the input signal can lead to significantly different coefficients or features for DFT and DWT.
Continuous wavelet transforms (CWT) with real-valued mother wavelets possess shift equivariance [13]. CWT coefficients offer a localized time-frequency representation of the signal, with scale-dependent locality [13]. This means they capture more localized features at higher frequencies and broader features at lower frequencies, while still maintaining overall locality compared to global transforms like DFT. While CWT has been used for TSC [15], global patterns of MTS data have been insufficiently considered. MTS data often contain sensible global patterns, such as trends, seasonality, periodicity, cycles, and long-term dependencies. Lacking the ability to capture shift-equivariant global features can lead to a loss of discriminative information for TSC.
These considerations highlight that many existing approaches cannot fully exploit the shift equivariance of features or patterns for TSC. While some CNN or CWT-based methods can capture shift-equivariant features, they often lack the ability to effectively utilize spectral or global features. A clear need arises for an effective approach that can leverage shift-equivariant local and global features in both time and frequency domains to enhance TSC performance. In this paper, we propose a novel approach for TSC that effectively leverages time shift-equivariant features and patterns from both time and frequency domains. The shift-equivariant spectral features are derived from the time-frequency representations of CWT with a real-valued mother wavelet. To remedy CWT’s limitation regarding the locality of resulting features, we leverage convolutional kernels to extract shift-equivariant temporal features.
While convolutional kernels excel at capturing temporal dependencies between input features or inter-dependencies between channels, preserving shift equivariance, they typically have limited lengths, resulting in a limited receptive field that extracts temporally localized content [16]. To enrich the expressiveness of CNN-based features, we adopt the kernel-based feature transformation ROCKET [4]. ROCKET uses random kernels with random lengths, potentially extracting a wide spectrum of temporal features, ranging from highly local to global ones.
However, many TSC tasks involve MTS data with characteristics or patterns spanning the entire length or a significant portion of the time series, suggesting that globally discriminative features or global interactions of features can be beneficial. While ROCKET may capture global features to a certain extent due to its use of kernels with random lengths, it tends to focus more on local features with varying degrees of locality. To enhance the extraction of global features, we incorporate fully connected MLPs. These can capture temporally global patterns with their wide receptive field, although they may not be well-suited for identifying temporally local patterns or temporal dependencies due to their treatment of each input feature independently. Thus, we use MLPs to complement ROCKET’s approach, strengthening the extraction of global features. Although hybrid approaches like ROCKET and MLP-based methods attempt to address local and global feature extraction, they do not inherently integrate spectral information or maintain shift equivariance across domains. Moreover, existing fusion mechanisms often lack adaptability, relying on fixed feature combinations rather than dynamic selection based on input characteristics. To address these limitations, we propose a novel feature fusion strategy that adaptively selects and integrates spectral and temporal features, leveraging their complementary strengths.
To avoid significantly increasing the size of extracted features, we propose a switch mechanism that selects between CNN-based (primarily local) features and MLP-based global patterns. This mechanism determines which type of temporal features - local or global - is more discriminative for a given input and integrates it with the CWT-domain features. Our experiments demonstrate that the switch mechanism effectively captures the most salient characteristics of the input data, whether they are local or global in nature.
The time-frequency representations from the CWT and temporal features from ROCKET transformation, or temporal features from global MLP, will be leveraged jointly to exploit the MTS characteristics, particularly shift equivariance, across domains. These different types of features provide various perspectives on the MTS data. We combine these perspectives with an approach known as multi-view learning, which has been shown to improve model performance and reliability, e.g., in image classification and clustering [17]. This provides comprehensive, enriched multiview contexts to the subsequent inference, thus enhancing TSC performance.
In addition, we introduce inversion invariance, a new concept in TSC where time series’ features or patterns are equally useful for classification when read in both forward and backward directions. This property is particularly relevant for data where the time direction is not inherently meaningful, such as ECG patterns, climate data, or rotational data with arbitrary start points. We posit that inversion invariance can generally enhance TSC performance for the following reasons:
  • 1.
    Using inversion invariance for TSC can effectively double the amount of input data, since both forward and backward readings will be used to train the same model. This increase in training examples is a new form of data augmentation, which leads to better generalization, reducing potential overfitting for TSC.
  • 2.
    Capturing inversion-invariant patterns may help enhance the model’s robustness to noise or disturbances. MTS data are often noisy [18], which can mask the underlying signals and affect the robustness of the algorithm. By identifying patterns meaningful in both directions, the TSC model may focus more on intrinsic patterns while being less affected by (potentially direction-specific) noise.
Inversion invariance can improve generalization and robustness for TSC, particularly when patterns are significant in both forward and reverse time directions. While Transformers and CNN-based architectures have been widely explored for time series modeling, their limitations in handling long sequences and computational overhead make them suboptimal for large-scale TSC tasks. SSMs like Mamba offer a promising alternative by efficiently modeling long-range dependencies with linear complexity. However, standard Mamba lacks explicit mechanisms for handling inversion invariance, a critical property in time-series data where temporal direction may be ambiguous (e.g., ECG signals). To overcome this, we introduce a novel scanning mechanism—tango scanning—that enhances inter-token relationships while preserving computational efficiency. Our extensive experiments across various MTS datasets demonstrate the effectiveness of this tango scanning approach, supported by theoretical analysis using attention mechanisms.
To facilitate final classification, we employ Mamba [19], a state-of-the-art (SOTA) model based on SSMs. Like RNNs, SSMs use state variables to represent the system’s internal condition and its evolution over time [20]. Mamba introduces selective state spaces, a mechanism that updates only a subset of state dimensions based on each input. This allows Mamba to focus on the most relevant information, efficiently process long sequences, and capture long-range dependencies. Mamba-based models have demonstrated competitive performance on various tasks, including language modeling [19, 21], time series forecasting [22], deoxyribonucleic acid (DNA) sequence modeling [19], tabular data learning [23], and audio generation [24]. Unlike popular Transformers, which have quadratic time complexity in sequence length, Mamba achieves linear time complexity. This makes it more suitable for processing long sequences and scaling to larger datasets. By using Mamba, our TSC model is efficient in training and inference, with reduced computational costs and memory requirements compared to existing SOTA models.
We introduce a novel sequence scanning scheme, tango scanning, for inputting the original sequence and the reversed sequence into the same Mamba block and fusing the output. This scheme uses essentially the same memory footprint but demonstrates higher accuracy than vanilla Mamba scanning. Our tango scanning scheme differs from existing bi-directional Mamba implementations in the following ways: 1) Compared to [25, 26], our approach uses one vanilla Mamba block for both directions, while theirs use two separate blocks. 2) Compared to BiMamba [27], we use a single Mamba block without weight ties, and only one reversal operation, whereas BiMamba uses two blocks with partial weight ties and two reversal operations. 3) Compared to MambaDNA [27], we use a single reversal operation and one Mamba block, while MambaDNA uses two “reverse complement” operations and two Mamba blocks with weight ties where the “reverse complement” operation is specific to DNA sequences to reflect the complementary nucleotides in the DNA base pairs. Through extensive experiments, we demonstrate that our multi-view Mamba-based approach outperforms or matches the performance of existing SOTA models, typically with a small fraction of computational requirements and reduced memory usage.
In summary, the contributions of this paper include:
  • 1.
@@
    Building a novel multi-view approach for TSC: Our approach seamlessly integrates frequency- and time-domain features to exploit shift-equivariant patterns and provide complementary, discriminative contexts for TSC. It also employs a gating scheme to fuse spectral features with local or global time-domain features, effectively leveraging patterns characterizing the MTS classes.
  • 2.
    Adapting the Mamba state-space model for sequence modeling and TSC: Our model leverages Mamba’s linear complexity to efficiently capture long-term dependencies within MTS data, while incorporating well-motivated modifications tailored for TSC tasks.
  • 3.
    Introduction of inversion invariance for TSC: This includes the new concept of inversion invariance for MTS classification. It also includes building an innovative Mamba-based “tango scanning” scheme to identify inversion invariant features or patterns. Our proposed tango scanning demonstrates improved effectiveness in modeling inter-token relationships in the sequence compared to vanilla Mamba block or bi-directional Mamba for TSC.
  • 4.
@@
    Extensive experimental validation of the proposed approach: Our model shows superior performance over various existing SOTA models over two sets of standard benchmarking datasets (10+20 datasets), achieving average accuracy improvements of 4.01-6.45% and 7.93% respectively over leading TSC models among 20 SOTA baselines.
By addressing key limitations of existing approaches—including inefficient long-range dependency modeling, suboptimal fusion of spectral and temporal features, and lack of inversion-invariant learning—our approach provides a well-rounded, efficient solution for TSC. The proposed TSCMamba framework not only achieves superior performance over SOTA baselines but also lays the groundwork for future research in multi-view learning and efficient sequence modeling for time series data. These contributions are expected to advance real-world TSC applications in research and everyday life. The following sections will briefly review related works, present our approach in detail, and demonstrate extensive experimental results and ablation studies to conclude the paper.
@@ 2.1 Traditional Methods
Traditional TSC methods include techniques like Dynamic Time Warping (DTW) [28], which measures the similarity between time series by aligning them in a non-linear way of dynamic programming. Tree-based methods like XGBoost [29] have also been applied to the TSC task.
2.2 Deep Learning Approaches
@@
In recent years, deep learning approaches have become increasingly popular for TSC. Various MLP-based methods have been proposed, including DLinear by Zeng et al. [30] and LightTS by Zhang et al. [31]. DLinear constructs a simple model based on MLP, while LightTS uses light sampling-oriented MLP. These models are generally efficient in computations. Convolutional neural networks (CNNs) have been adapted for TSC, such as ROCKET [4, 32] which uses random convolutional kernels for fast and accurate classification. ROCKET has inspired multiple extensions, including MiniROCKET [33] and MultiROCKET [34], which further optimize feature extraction efficiency and classification performance. These methods leverage thousands of randomly initialized convolutional kernels to capture diverse temporal patterns without requiring extensive training, making them effective for time series classification. Given its strength, ROCKET-based features have been widely integrated into hybrid models. Inspired by this, TSCMamba leverages ROCKET-derived representations to complement spectral features, enhancing discriminability in time series classification. CNN has been also used by Franceschi et al. [5] for learning representations of multivariate time series in an unsupervised way, which is then further leveraged for TSC. Besides CNNs, recurrent neural networks (RNNs) such as long-short-term memory (LSTM) [35] and a variant, a gated recurrent unit (GRU), has also been adopted for TSC. Moreover, CNN and RNN have been combined to handle TSC effectively [36].
2.3 Transformer-Based Methods
@@
Transformers by Vaswani et al. [37], originally used for natural language processing, have been adapted for time series modeling. Reformer by Kitaev et al. [38] introduces efficiency improvements to handle longer sequences. Numerous Transformer variants have been proposed to better model the unique characteristics of time series, such as handling non-stationarity with on-stationary Transformers by Liu et al. [39], combining exponential smoothing with ETSformer in Woo et al. [40], and using decomposition and auto-correlation with Autoformer in Wu et al. [41]. Other variants include Pyraformer [42], which uses pyramidal attention to reduce complexity, Flowformer [43], which linearizes Transformers using conservation flows, and Informer [44], which focuses on efficient long sequence forecasting by exploiting frequency-enhanced decomposition. Notably, TimesNet [45] models the temporal 2D variations in time series data using a hierarchical structure of temporal blocks. By combining 2D convolutions, multi-head self-attention, and a novel positional encoding scheme, it can capture both local patterns and long-range dependencies in time series and obtain state-of-the-art performance. Recently, patch-based methods have gained attention in time series modeling, particularly in Transformer-based architectures. PatchTST [46] introduces a non-overlapping patching mechanism to improve long-range dependency modeling. Unlike sequence-to-sequence methods, PatchTST processes time series data in local patches, capturing finer-grained temporal structures. Despite the impressive performance of Transformer-based models, Zeng et al. [30] have shown that MLP-based models can be more effective in many scenarios.
2.4 State-Space Models
Recently, a method called LSSL by Gu et al. [47] has been proposed for TSC with a structured SSM called S4. It employs a special parameterization called the Diagonal Plus Low-Rank form to represent the state transition matrix, enabling efficient computation over long sequences. Mamba [19] dynamically parameterizes state transitions based on input content, filtering out irrelevant information and selectively propagating key features. This mechanism has demonstrated strong results in sequence modeling, motivating its adaptation in TSCMamba to enhance time series classification performance. These innovations have made SSMs competitive alternatives to traditional RNN and Transformer-based models. Furthermore, recent works such as TimeMachine [22] and S-Mamba [25] have demonstrated the potential of Mamba in time series forecasting, reinforcing the effectiveness of structured state-space models in capturing long-range dependencies.